OpenStudy (anonymous):

A manufacturer received an order of 210 computer chips. Unfortunately, 10 of the chips are defected. To test the shipment, a quality-control engineer randomly selects 24 chips from the box of 210 and tests them. The random variable X represents the number of defective chips in the sample. What is the probability of obtaining 5 defective chips? What is the probability of obtaining 3 defective chips? What is the probability that the QC engineer will not find any defective chips? Probability of finding 13 defective chips? How many defective chips would you expect to select?

OpenStudy (anonymous):

@jim_thompson5910 last one i need help with!

jimthompson5910 (jim_thompson5910):

well right off the bat, I can answer "Probability of finding 13 defective chips?" without having to do any math at all really do you see what I mean? or no?

OpenStudy (anonymous):

yeah that one is 0 right?

OpenStudy (anonymous):

since theres only 12?

jimthompson5910 (jim_thompson5910):

good, it's impossible since there are only 10 defective chips

OpenStudy (anonymous):

or 10 i mean

OpenStudy (anonymous):

i've tried everything to find the others and im stumped :/

jimthompson5910 (jim_thompson5910):

let me think on those

OpenStudy (anonymous):

no problem!

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (anonymous):

TI-84

jimthompson5910 (jim_thompson5910):

ok we'll have to rely on the nCr function do you know how to find that?

OpenStudy (anonymous):

yes!

jimthompson5910 (jim_thompson5910):

have you learned about the hypergeometric distribution?

OpenStudy (anonymous):

no :/

jimthompson5910 (jim_thompson5910):

That distribution relies on the nCr formula

jimthompson5910 (jim_thompson5910):

we pick a sample of 24 chips

jimthompson5910 (jim_thompson5910):

in this sample, we want to know the probability of having exactly 5 bad chips

jimthompson5910 (jim_thompson5910):

There are 10 bad chips overall So there are 10 C 5 ways to pick 5 bad chips. We won't calculate 10 C 5 since we'll let the calculator do this for us

OpenStudy (anonymous):

so i got 252?

jimthompson5910 (jim_thompson5910):

our sample size is 24 if we allot 5 spaces for the bad chips, we have 24 - 5 = 19 spaces for good chips

jimthompson5910 (jim_thompson5910):

There are 210 - 10 = 200 good chips so there are 200 C 19 ways to pick 19 good chips (from a pool of 200) if you were to calculate 200 C 19 you'll get a very large number

jimthompson5910 (jim_thompson5910):

put together, there are (10 C 5)*(200 C 19) ways to pick 5 bad chips AND 19 good chips

jimthompson5910 (jim_thompson5910):

this is out of 214 C 24 ways to pick 24 chips total

jimthompson5910 (jim_thompson5910):

so in your calculator, you will type this in ( (10 nCr 5)*(200 nCr 19) )/( 214 nCr 24 )

OpenStudy (anonymous):

i got .00125

jimthompson5910 (jim_thompson5910):

oh sry, not 214, 210

jimthompson5910 (jim_thompson5910):

fixed ( (10 nCr 5)*(200 nCr 19) )/( 210 nCr 24 )

OpenStudy (anonymous):

oh okay now i got .002019

jimthompson5910 (jim_thompson5910):

here's what I get

jimthompson5910 (jim_thompson5910):

and you got the same

OpenStudy (anonymous):

is the next the same just substitute 3 for 5?

jimthompson5910 (jim_thompson5910):

for the next part, you are changing the "5 bad chips" to "3 bad chips" so you'll have ( (10 nCr 3)*(200 nCr 21) )/( 210 nCr 24 )

jimthompson5910 (jim_thompson5910):

the 5 changed to 3 the 19 changed to 21

jimthompson5910 (jim_thompson5910):

also I recommend you read this page http://staffwww.fullcoll.edu/kchang/MATH%20120%20Primavera%202012/6.4_HypergeometricProbability.pdf as it has some other examples on the hypergeometric distribution

OpenStudy (anonymous):

.0746?

jimthompson5910 (jim_thompson5910):

getting the same

OpenStudy (anonymous):

thanks i will!

jimthompson5910 (jim_thompson5910):

for part c, you'll have 0 defective chips in the sample so you'll have ( (10 nCr 0)*(200 nCr 24) )/( 210 nCr 24 )

OpenStudy (anonymous):

.28875?

jimthompson5910 (jim_thompson5910):

same

jimthompson5910 (jim_thompson5910):

part d is already done

OpenStudy (anonymous):

sweet! just one more :)

jimthompson5910 (jim_thompson5910):

part e will require the formula in the blue box on page 5 http://staffwww.fullcoll.edu/kchang/MATH%20120%20Primavera%202012/6.4_HypergeometricProbability.pdf

jimthompson5910 (jim_thompson5910):

the first formula in that box

OpenStudy (anonymous):

okay i see it

jimthompson5910 (jim_thompson5910):

n = 24 (sample size) k = 10 (number of defective chips...ie number of "successes") N = 210 (overall population size)

OpenStudy (anonymous):

.9842? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

use the n*k/N formula

OpenStudy (anonymous):

ah i was hoping i got it right after all that

OpenStudy (anonymous):

oh 1.14?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

n*k/N = 24*10/210 = 240/210 = 1.142857 so you are correct

OpenStudy (anonymous):

thank you so much Jim! i know it was a lot but i appreciate it so much

jimthompson5910 (jim_thompson5910):

sure thing